OpenStudy (anonymous):

HELP STEP BY STEP INTEGRATION (I mean step by step!!)\[\int\limits {x \over \sqrt {x-3}}dx--->x=3+u^2\]

OpenStudy (experimentx):

simplify .. first!!

OpenStudy (anonymous):

6u + (2/3)u^3

OpenStudy (anonymous):

\[\int\limits {3+u^2 \over \sqrt{3+u^2-3}}\]

OpenStudy (anonymous):

\[\int\limits\limits{ 3+u^2 \over u} \times 2u\]

OpenStudy (anonymous):

\[2 \int\limits {3u+u^3 \over u}\]

OpenStudy (anonymous):

How to go on from here?

OpenStudy (experimentx):